OpenStudy (anonymous):

Evaluate (ab)^2 for a = 4 and b = 3. A. 36 B. 24 C. 81 D. 144

OpenStudy (anonymous):

12*12

OpenStudy (lgbasallote):

\[\LARGE [(4)(3)]^2\]

OpenStudy (anonymous):

Simple: (ab)^2 = (4-3)^2 = 12^2 (not 12^12) = 144

OpenStudy (lgbasallote):

@dpaInc said 12*12 (12 x 12) which also means 12^2 lol

OpenStudy (anonymous):

OOPS

OpenStudy (anonymous):

I'm very sleepy, can't see properly ...

OpenStudy (lgbasallote):

12*12 (12 x 12) does not mean multiplication lol...the one in parenthesis is another interpretation